OpenStudy (anonymous):

For the equation find the equation of the graph that is parallel and perpendicular that passes through the specified point. Then graph the equations 3x - y = 9

OpenStudy (karatechopper):

ok so basically lets start off with what the parallel slope is...so whats the parallel slope of 3x-y=9 @MSFFL

OpenStudy (anonymous):

-y = -3x + 9 which ends up being y = 3x +9 so 3 is the slope

OpenStudy (campbell_st):

what is the specified point

OpenStudy (anonymous):

the specified points are (1,1)

OpenStudy (karatechopper):

now how does -y=-3x+9 become y=3x+9? impossible

OpenStudy (anonymous):

two signs must change for the problem to remain the same

OpenStudy (anonymous):

I learned it early on when I was changing sign in fractions

OpenStudy (campbell_st):

the parallel line has the same gradient the given line in y = 3x - 9 m = 3 so the parallel line is y - 1=3(x - 1) y =3x -2 the perpendicular line has a gradient of -1/m in this question its -1/3 so y - 1 = -1/3(x -1) y = 1/3x + 4/3

OpenStudy (anonymous):

I got that far now graphing is what I am having problems with how to you graph 1. y = 3x +9 2. y = 3x - 2 and 3. y = -1/3 + 4/3

OpenStudy (karatechopper):

ok so remember y=mx+b form? b=y intercept

OpenStudy (anonymous):

yep

OpenStudy (campbell_st):

just use a table of values for the line y = 3x -2